# Human Reactions to Incorrect Answers from Robots

### v1

@@ Abstract
As robots grow more and more integrated into numerous industries, it is critical to comprehend how humans respond to their failures. This paper systematically studies how trust dynamics and system design are affected by human responses to robot failures. The three-stage survey used in the study provides a thorough understanding of human-robot interactions. While the second stage concentrates on interaction details, such as robot precision and error acknowledgment, the first stage collects demographic data and initial levels of trust. In the last phase, participants’ perceptions are examined after the encounter, and trust dynamics, forgiveness, and propensity to suggest robotic technologies are evaluated. Results show that participants’ trust in robotic technologies increased significantly when robots acknowledged their errors or limitations to participants and their willingness to suggest robots for activities in the future points to a favorable change in perception, emphasizing the role that direct engagement has in influencing trust dynamics. By providing useful advice for creating more sympathetic, responsive, and reliable robotic systems, the study advances the science of human-robot interaction and promotes a wider adoption of robotic technologies.
@@ I Introduction
Figure 1: Participants are interacting with the robot
Several recent studies have shed light on various facets of human-robot interaction and responses to robot errors. In [5], the researchers highlight the significance of multimodal cues in detecting and addressing conversational failures in HRI. This research underscores the complexity of human responses to errors in robot communication, emphasizing the role of non-verbal cues in this process.
@@
    Humans are more forgiving of informational errors when they come from familiar sources like other humans or internet search engines, compared to robots.
This study is conducted through a three-stage survey designed to capture a comprehensive view of human-robot interactions. By systematically investigating these aspects, this research aims to bridge a critical knowledge gap in HRI. It contributes to a deeper understanding of human-robot trust dynamics and provides insights for developing more effective, reliable, and user-friendly robotic systems. This study is poised to enhance the field of HRI by offering a nuanced perspective on how human reactions to robotic errors can inform the design of more empathetic, responsive, and context-aware robots.
@@ II Related Work
III Experimental Setup
System Configuration: In this experiment, we used the NAO robot, a humanoid platform renowned for its agility and interactive capabilities depicts in Fig. 3. Nao was positioned directly across from the participants in a lab environment. It started the session by greeting them and encouraging them to ask questions from a predetermined set of ten questions designed to evaluate its conversational abilities. Nao is equipped with advanced sensors and a 1.91 GHz Quad core Atom E3845 Processor and 4 GB DDR3 ram , which enables real-time data collection and efficient processing of complex interactions. This combination of features facilitates a comprehensive assessment of human-robot dialogue. It also highlights Nao’s suitability for intricate interaction studies, allowing it to navigate the nuances of human engagement effectively through computational intelligence.
Hardware and Software Integration: The Nao robot was interconnected with a personal computer serving as the speech recognition and response generation processing unit. This integration was facilitated by the Robot Operating System (ROS), which provided a flexible framework for controlling the robot. The speech interaction mechanism was engineered such that the participant’s spoken queries were captured by Nao’s built-in microphone, transmitted to the computer, and processed using an advanced speech recognition application.
Speech Recognition and Natural Language Processing: Our speech recognition application leveraged the ”speech_recognition” Python library to convert spoken language into text form. Leveraging the Natural Language Toolkit (NLTK), we performed various natural language processing (NLP) tasks such as tokenization, lemmatization, and stop-word analysis. These NLP techniques enabled the robot to comprehend human language more effectively by allowing it to discern the context and significance of the queries posed by the participants.
@@ III Experimental Setup
Implementation of Error Handling and Apology Mechanisms: To investigate the impact of robot errors on human trust and perception, our experimental protocol deliberately included a mechanism for generating incorrect responses. A randomized algorithm intermittently selected responses that were intentionally erroneous, simulating real-world scenarios where robots may misinterpret or incorrectly process human input. Upon issuing a mistaken response, our system was designed to automatically detect the anomaly and initiate an apology sequence. This sequence involved the Nao robot acknowledging its error and expressing regret to the participant, either through a direct apology or, in some cases, by attempting to correct the misinformation with the accurate answer. This approach not only allowed us to observe participant reactions to robot failures but also provided insights into the dynamics of trust recovery and forgiveness in human-robot interactions. By integrating these error-handling and apology mechanisms, our study offers a comprehensive view of how such factors influence the overall efficacy and acceptance of robotic technologies in interactive settings.
@@ IV Methodology
The research project’s methodology is meticulously planned to gauge participants’ initial emotional responses and reflective thoughts following interactions with a robot. A mixed-methods strategy is applied, integrating qualitative and quantitative data via a three stage survey approach.
IV-A Three Stage Study
@@ IV-A1 Preliminary Survey
This stage collects demographic data and baseline attitudes toward robots. We conducted this survey before the robot interaction which involves a detailed survey with specific questions designed to gather background information and preconceptions about robots and automation from the participants. The survey includes Age Group, Gender, Country, Prior Interaction with Robots, Perception of Robots, and Trust in Automation. This survey sets a baseline for understanding each participant’s initial stance towards robots and automated systems, which is essential for analyzing any shifts in perception following the robot interaction.
IV-A2 Interaction with Robot
@@
IV-A3 Reflective Survey
Assesses the change in perceptions and trust levels post-interaction. This stage is crucial in the methodology for gathering vital survey data post-interaction with Nao. This stage delves into the participants’ nuanced reactions and perceptions, focusing on several key areas such as Error Impact on Trust, Trust Impact of Unacknowledged Error, Comparative Forgiveness (Human vs. Robot), Effectiveness of Robot’s Apology, Importance of Acknowledging Errors, Willingness to Rely on Robot Post-Error, Future Interaction Intent, Perceived Reliability of Robots, and Recommendation Likelihood. This stage is integral to the study as it seeks to understand the depth and breadth of human responses to robotic performance, particularly in the context of trust and reliability. The collected data will provide rich insights into the dynamics of human-robot interaction and the psychological impacts of robotic errors and their acknowledgments.
The methodology depicted in Fig. 2 maps out the process of a human-robot interaction study. Initially, participants complete a survey gauging their demographics, education, prior experience with robots, and trust in automation. They then interact with NAO[12], in a series of interactions during which they pose predetermined questions as shown in the Fig. 1. Following the interaction, participants complete two further surveys that capture their immediate and reflective responses to the robot’s performance. The collected data from these surveys is methodically stored for in-depth analysis to conclude the impact of robotic interactions on human trust and perception. The design allows for a comprehensive analysis of the human-robot interface, capturing the depth and evolution of participant responses.
@@ IV-B Participant Selection
The participant group for this study comprises 31 individuals from different countries between 21 and 64 years of age where 9 of them are female and 22 of them are male. This group is largely inexperienced in interacting with robots previously, as most participants(17 out of 31) are engaging in such interactions for the first time. Their general perception of robots leans towards viewing them as ”somewhat reliable and useful.” The trust in automation technology among these participants varies, with a moderate level of trust being the most common trend. These insights suggest that the group, while diverse, shares an optimism about the potential of robotic systems despite limited prior exposure. This demographic appears open to advancements in human-robot interactions (HRI), emphasizing the significance of building trust in automation for broader acceptance of future technologies. The findings point towards the necessity of developing user-friendly interfaces and ensuring transparency in robot functionalities, which could further improve perceptions and trust in such technologies.
IV-C Interaction with Robot
@@
In the robot interaction stage of the study, participants individually engage with the Nao, a programmatically error-prone robot. During this phase, they are presented with a predefined list of 10 general knowledge questions to ask the Nao. These questions cover topics such as historical dates, geographical facts, and common knowledge queries such as the independence day of the US, capital of France, the first US president, shape of a stop sign, and so on. Nao, equipped with a programmed response set, deliberately answers some questions incorrectly. Notably, after each incorrect response, Nao is designed to acknowledge its mistake and issue an apology to the participant. This interaction phase is pivotal for observing and analyzing the participants’ real-time reactions to robotic errors and apologies, which is crucial for understanding human-robot interaction dynamics. During this interaction, Nao is programmed to answer some questions incorrectly and then acknowledge and apologize for these errors, allowing for observation of participant reactions to the robot’s mistakes and subsequent apologies.
@@ V Results and Discussion
  • Perceived Reliability and Usefulness: Participants with limited prior interaction with robots perceived them as somewhat reliable and useful. The variation in trust levels among participants indicates a nuanced perspective, about 58% of participants expressed moderate trust in robots before the interactions with the robot.
@@
  • Reaction to Errors: All of the participants observed errors in robot responses. Interestingly, about 84% of participants reported feeling more trusting when the errors were acknowledged and apologized for, highlighting the potential impact of transparent communication on user trust.
  • Willingness to Rely on Robots: Despite noticing errors, many participants remained willing to rely on robots for future tasks. This suggests a recognition among users that robots, like humans, possess both strengths and weaknesses.
  • Trust Post-Mistakes: Participants’ trust in robots either remained unchanged or increased after witnessing mistakes. This finding suggests a growing acceptance of robots, with many participants open to considering their use in various domains.
Figure 4: Perception on Robots Before and After the Study
To support the first hypothesis (H1), asserting that ”The extent to which a robot acknowledges its limitations or errors can significantly enhance human trust and emotional acceptance” participants were asked to compare the forgiveness between human and robot. The result shows that humans are more forgiving of informational errors when they come from familiar sources like other humans or internet search engines, compared to robots. The data reveals that more than 25% of participants indicated they would forgive the robot more than the humans, while approximately 55% were less forgiving to the robot. Fig. 4 shows the perception on robots, before and after the study. So that we can say our first hypothesis is supported.
To support the second hypothesis (H2), we compared the trust level on the robot before and after the study. Experimental result shows that about 32% of participants had low trust in automation before the experiment but after the experiment, the trust level increased to around 84% and around 84% of participants were likely to recommend the use of robots in high-stakes domains. Fig. 5 shows the visual representation of these statistics. So that we can say our second hypothesis is supported.
Figure 5: People are more forgiving to human than robot

### v2

# Human-Robot Interaction and Perceived Irrationality: A Study of Trust Dynamics and Error Acknowledgment

@@ Abstract
As robots become increasingly integrated into various industries, understanding how humans respond to robotic failures is critical. This study systematically examines trust dynamics and system design by analyzing human reactions to robot failures. We conducted a four-stage survey to explore how trust evolves throughout human-robot interactions. The first stage collected demographic data and initial trust levels. The second stage focused on preliminary expectations and perceptions of robotic capabilities. The third stage examined interaction details, including robot precision and error acknowledgment. Finally, the fourth stage assessed post-interaction perceptions, evaluating trust dynamics, forgiveness, and willingness to recommend robotic technologies. Results indicate that trust in robotic systems significantly increased when robots acknowledged their errors or limitations. Additionally, participants showed greater willingness to suggest robots for future tasks, highlighting the importance of direct engagement in shaping trust dynamics. These findings provide valuable insights for designing more transparent, responsive, and trustworthy robotic systems. By enhancing our understanding of human-robot interaction (HRI), this study contributes to the development of robotic technologies that foster greater public acceptance and adoption.
@@ I Introduction
Figure 1: Participants interact with the robot
Recent studies have shed light on various facets of human-robot interaction and responses to robot errors. In [5], the researchers highlight the significance of multimodal cues in detecting and addressing conversational failures in HRI. This research underscores the complexity of human responses to errors in robot communication, emphasizing the role of non-verbal cues in this process.
@@
    Compared to robots, humans are more forgiving of informational errors when they come from familiar sources like other humans or internet search engines.
This study is conducted through a four-stage survey designed to capture a comprehensive view of human-robot interactions. By systematically investigating these aspects, this research aims to bridge a critical knowledge gap in HRI. It contributes to a deeper understanding of human-robot trust dynamics and provides insights for developing more effective, reliable, and user-friendly robotic systems. This study is poised to enhance the field of HRI by offering a nuanced perspective on how human reactions to robotic errors can inform the design of more empathetic, responsive, and context-aware robots.
@@ II Related Work
III Experimental Setup
System Configuration: In this experiment, we used the NAO robot[12], a humanoid platform renowned for its agility and interactive capabilities, as depicted in Fig. 3. The NAO robot was selected for this study due to its advanced capabilities, adaptability, and widespread use in human-robot interaction (HRI) research. NAO is designed to facilitate naturalistic and interactive conversations, making it an ideal platform for studying trust dynamics in human-robot interactions.
One of the primary reasons for choosing NAO is its rich set of interactive features, including speech recognition, natural language processing, and expressive gestures. These capabilities allow NAO to engage with participants in a way that mimics human-like conversational behavior, making it well-suited for evaluating how humans respond to robotic errors.
NAO was positioned directly across from the participants in a controlled lab environment. It initiated the session by greeting them and encouraging them to ask questions from a predetermined set of ten questions designed to evaluate its conversational abilities. NAO is equipped with advanced sensors, a 1.91 GHz Quad-Core Atom E3845 Processor, and 4 GB DDR3 RAM, which enable real-time data collection and efficient processing of complex interactions. This combination of features facilitates a comprehensive assessment of human-robot dialogue while highlighting NAO’s suitability for intricate interaction studies. Its computational intelligence allows it to navigate the nuances of human engagement effectively.
Furthermore, NAO has been widely adopted in HRI research, education, and cognitive science due to its human-like form factor and versatility in experimental settings. Its ability to maintain eye contact, use body language, and recognize user inputs makes it particularly useful for studying trust-building mechanisms in robots. NAO provides an optimal balance between computational power, real-time interaction capabilities, and human-like engagement, making it the preferred choice for our study.
@@ III Experimental Setup
Hardware and Software Integration: The NAO robot was integrated with a personal computer that functioned as the primary speech recognition and response generation processing unit. This setup leveraged the Robot Operating System (ROS), which provided a modular and extensible framework for robot control, real-time data processing, and inter-device communication. The integration ensured efficient handling of computationally intensive tasks while allowing for seamless synchronization between NAO’s hardware components and external processing units.
The speech interaction mechanism was designed to facilitate real-time, high-accuracy query processing. Spoken queries from participants were captured by NAO’s built-in microphone array, which supports multi-directional audio input for enhanced speech detection. The audio signals were then transmitted to the external computing unit via a low-latency communication protocol, where they were processed using an Automatic Speech Recognition (ASR) engine optimized for natural language understanding (NLU). The recognized text was subsequently passed through a Natural Language Processing (NLP) pipeline, which incorporated error correction, intent recognition, and contextual analysis to generate coherent responses. The final responses were synthesized using text-to-speech (TTS) processing, enabling NAO to deliver verbal output in a human-like manner.
Additionally, the integration enabled real-time monitoring and logging of participant interactions, allowing for data collection, system debugging, and performance analysis. By leveraging ROS middleware, external processing resources, and AI-driven speech models, this setup provided a robust, scalable, and adaptive environment for evaluating human-robot conversational dynamics in controlled experimental settings. Fig. 4 illustrates the hardware-software architecture, showcasing the interaction flow between NAO, the processing unit, and the speech recognition system.
Figure 4: Hardware and Software Integration of the NAO Robot
Speech Recognition and Natural Language Processing: Our speech recognition application leveraged the ”speech-recognition” Python library to convert spoken language into text form. Leveraging the Natural Language Toolkit (NLTK), we performed various natural language processing (NLP) tasks such as tokenization, lemmatization, and stop-word analysis. These NLP techniques enabled the robot to comprehend human language more effectively by allowing it to discern the context and significance of the queries posed by the participants.
@@
Implementation of Error Handling and Apology Mechanisms: To investigate the impact of robot errors on human trust and perception, our experimental protocol deliberately included a mechanism for generating incorrect responses. A randomized algorithm intermittently selected responses that were intentionally erroneous, simulating real-world scenarios where robots may misinterpret or incorrectly process human input. Upon issuing a mistaken response, our system was designed to automatically detect the anomaly and initiate an apology sequence. This sequence involved the Nao robot acknowledging its error and expressing regret to the participant, either through a direct apology or, in some cases, by attempting to correct the misinformation with an accurate answer. This approach not only allowed us to observe participant reactions to robot failures but also provided insights into the dynamics of trust recovery and forgiveness in human-robot interactions. By integrating these error-handling and apology mechanisms, our study offers a comprehensive view of how such factors influence the overall efficacy and acceptance of robotic technologies in interactive settings.
@@ IV Methodology
The research project’s methodology is meticulously planned to gauge participants’ initial emotional responses and reflective thoughts following interactions with a robot. A mixed-methods strategy is applied, integrating qualitative and quantitative data via a four-stage survey approach.
IV-A Four Stage Study
@@ IV-A1 Preliminary Survey
This stage collects demographic data and baseline attitudes toward humans as well as robots. We conducted this survey before the human and robot interaction which involves a detailed survey with specific questions designed to gather background information and preconceptions about robots and automation from the participants. The survey includes Age Group, Gender, Country, Prior Interaction with Robots, Perception of Robots, and Trust in Automation. This survey sets a baseline for understanding each participant’s initial stance towards humans and automated systems, which is essential for analyzing any shifts in perception following the human and robot interaction.
IV-A2 Perception of Interaction with Humans
This stage occurs after the participants’ interaction with a human. In this stage, participants answer a series of survey questions about their experience. They are asked about the number of questions they ask the human, evaluate the correctness of the human’s responses, indicate their awareness of any errors made by the human, and observe whether the human apologized for its mistakes. This post-interaction survey is crucial for capturing the participants’ immediate reactions and perceptions of the human’s accuracy and response to errors. This stage is designed to compare reactions between human-to-human and human-to-robot interactions.
IV-A3 Perception of Interaction with Robot
@@
IV-A4 Reflective Survey
Assesses the change in perceptions and trust levels post-interaction. This stage is crucial in the methodology for gathering vital survey data post-interaction with the Human and Nao. It delves into the participants’ nuanced reactions and perceptions, focusing on several key areas such as the Impact of Error on Trust, Trust Impact of Unacknowledged Errors, Comparative Forgiveness (Human vs. Robot), Effectiveness of Robot’s Apology, Importance of Acknowledging Errors, Willingness to Rely on Robot Post-Error, Future Interaction Intent, Perceived Reliability of Robots, and Likelihood of Recommendation. This stage is integral to the study as it seeks to understand the depth and breadth of human responses to robotic performance, particularly in the context of trust and reliability. The collected data will provide rich insights into the dynamics of human-robot interaction and the psychological impacts of robotic errors and their acknowledgments.
The methodology depicted in Fig. 2 maps out the process of human-robot and human-human interaction studies. In the first stage, after completing the preliminary survey, humans interact with a human, engaging in a series of interactions during which they pose predetermined questions. In stage two, after interacting with the human, humans interact with NAO. Here, humans also engage in a series of interactions during which they pose predetermined questions. Following the interaction, participants complete two further surveys that capture their immediate and reflective responses to the human’s and robot’s performance. The collected data from these surveys are methodically stored for in-depth analysis to conclude the impact of robotic interactions on human trust and perception. The design allows for a comprehensive analysis of the human-robot and human-human interface, capturing the depth and evolution of participant responses.
@@ IV-B Participant Selection
The participant group for this study consists of 31 individuals from various countries, aged between 21 and 64 years old, with 9 females and 22 males. Most participants (17 out of 31) have little to no prior experience in interacting with robots, making this study their first significant exposure to such technology (Figure 5). Generally, their perception of robots tends to lean towards viewing them as ’somewhat reliable and useful.’ However, there is variability in their trust levels towards automation technology, with a moderate level of trust being the most common trend observed. These insights indicate that while the participant group is diverse, there exists an overall optimism about the potential of robotic systems despite limited prior exposure. Moreover, they demonstrate openness to advancements in human-robot interactions (HRI), highlighting the importance of establishing trust in automation for broader acceptance of future technologies. The findings underscore the necessity of developing user-friendly interfaces and ensuring transparency in robot functionalities, as these factors could significantly enhance perceptions and trust in such technologies.
Figure 5: Participant Selection
IV-C Questions Selection
To evaluate the robot’s ability to respond to factual inquiries, a set of ten general knowledge questions was selected. These questions were designed to cover a broad range of topics, including history, geography, science, mathematics, language, and symbolic recognition. Each question requires a definitive, objective answer, allowing for a straightforward assessment of the robot’s accuracy.
The selected questions are as follows:
  1. 1.
    What is the date of Independence Day in the USA?
  2. 2.
    How many stars are on the flag of the USA?
  3. 3.
    What is the capital of France?
  4. 4.
    What is the name of the largest planet in our solar system?
  5. 5.
    What is the currency of the USA?
  6. 6.
    How many sides does a triangle have?
  7. 7.
    What is the opposite of ‘day’?
  8. 8.
    What shape is a ‘stop’ sign?
  9. 9.
    Who was the first president of the USA?
  10. 10.
    Who wrote the play Romeo & Juliet?
These questions serve multiple purposes. Firstly, they assess the robot’s factual recall capabilities, ensuring it can provide accurate, well-structured responses to commonly known information. Secondly, questions such as #8 (stop sign shape) test the robot’s ability to recognize real-world symbols, which is crucial for navigation and contextual understanding in human-robot interactions. Additionally, simple linguistic relationships are examined, as seen in #7 (opposite of ‘day’), allowing for an assessment of the robot’s language processing ability.
While these questions focus on fact-based knowledge, future work may expand the scope to include multi-step reasoning tasks and handling of ambiguous questions. This would provide a more comprehensive evaluation of the robot’s cognitive processing and adaptability in complex conversational settings.
IV-D Interaction with Human
In the human interaction stage of the study, participants individually engage with a human. During this stage, they are presented with a predefined list of 10 general knowledge questions to ask the human. We designed this stage so that the human intentionally provides some random incorrect answers. After each incorrect answer, the human acknowledges the mistake and apologizes to the participants.
IV-E Interaction with Robot
@@
In the robot interaction stage of the study, participants individually engage with the Nao, a programmatically error-prone robot. During this phase, they are presented with the same predefined list of 10 general knowledge questions to ask the Nao. Nao, equipped with a programmed response set, deliberately answers some questions incorrectly. Notably, after each incorrect response, Nao is designed to acknowledge its mistake and issue an apology to the participant. This interaction phase is pivotal for observing and analyzing the participants’ real-time reactions to robotic errors and apologies, which is crucial for understanding human-robot interaction dynamics. During this interaction, Nao is programmed to answer some questions incorrectly and then acknowledge and apologize for these errors, allowing for observation of participant reactions to the robot’s mistakes and subsequent apologies.
@@ V-A Trust and Forgiveness in Human-Robot Interaction
  • Perceived Reliability and Usefulness: Participants with limited prior experience with robots generally found them somewhat reliable and useful. Trust levels varied significantly, with 58% expressing moderate trust before interacting with the robot. However, some participants remained skeptical, suggesting that initial trust in automation is influenced by personal experience and exposure.
  • Reaction to Errors: All participants noticed errors in the robot’s responses. However, 84% reported increased trust when the robot acknowledged and apologized for its mistakes. This underscores the importance of transparency in human-robot interaction, as error acknowledgment appears to enhance trust rather than diminish it.
  • Willingness to Rely on Robots: Despite observing mistakes, many participants expressed confidence in using robots for future tasks. This suggests a growing acceptance of robots, where users recognize that even intelligent systems can make errors just like humans. Designing robots that communicate their limitations effectively could further strengthen their perceived reliability.
  • Trust Post-Mistakes: After witnessing errors, trust in the robot either remained unchanged or increased for most participants. This indicates that error-handling strategies (such as apologies and explanations) play a crucial role in trust dynamics. The results suggest that when robots acknowledge their errors, users are more likely to maintain or even increase their trust in automation.
  • Forgiveness of Human Errors: Humans are generally more forgiving of mistakes made by other humans compared to those made by robots. This is largely due to social and cognitive biases, where human errors are often attributed to situational factors such as distraction, fatigue, or momentary lapses in attention. In contrast, robotic errors are more likely to be perceived as system failures, raising concerns about reliability and competence. This tendency aligns with our findings, where participants were significantly less forgiving toward robots than humans or search engines. These results suggest that if robots incorporate more human-like transparency, social cues, and adaptive learning behaviors, they may be perceived as more trustworthy and acceptable in real-world applications.
Figure 6: Perception of Robots Before and After the Study
Figure 7: People are more forgiving to humans than robots
V-B Evaluation of Hypotheses
V-B1 Support for H1: The impact of error acknowledgment on trust
@@
To evaluate H1, all participants were exposed to errors in the robot’s responses. Interestingly, 84% of participants reported increased trust when the robot acknowledged its errors and apologized, highlighting the importance of transparent communication in fostering user trust.
A comparative analysis of participants’ trust levels before and after the study further reinforces this finding. As shown in Figure 6, only 32.3% of participants initially perceived the robot as reliable, with 19.4% considering it unreliable and 48.4% remaining neutral. However, after interacting with the robot, 83.9% of participants reported increased reliability perception, while only 6.5% experienced a decrease in trust. Additionally,84% of participants expressed a willingness to recommend the use of robots in high-stakes domains.
These results strongly support H1, demonstrating that when robots acknowledge their limitations and errors, human trust and emotional acceptance significantly increase.
V-B2 Support for H2: Forgiveness of errors from different sources
To assess H2, participants were asked to compare their level of forgiveness toward informational errors made by humans, robots, and internet search engines. The results indicate that humans are generally more forgiving of errors from familiar sources compared to robots.
As illustrated in Figure 7, 54.8% of participants were less forgiving toward robots, whereas 25.8% were more forgiving toward robots than humans, and 19.4% expressed equal levels of forgiveness for both sources. These findings suggest that people hold robots to a higher standard of accuracy than human or internet-based sources, possibly due to differing expectations regarding machine intelligence and perceived competence.
These findings provide strong empirical support for H2, confirming that humans are more likely to overlook errors when they come from familiar sources rather than from robots.
V-C Statistical Analysis of Trust Level Changes